OpenStudy (anonymous):

pipe A can fill tank in 3 hours.If pipe B can fill the same tank 2 hours,how many minutes will it take both pipes to fill 2/3 of the tank?

OpenStudy (anonymous):

let volume of tank = x rate of A = x/3 , rate of B = x/2 time taken = y (hours) y.x/3 + y.x/2 = 2/3 * x yx(1/3 + 1/2) = 2/3*x X CANCELS y(2/6 + 3/6) = 2/3 y(5/6) = 2/3 [ 6 & 3 cancel] 5y = 2*2 y = 4/5 hours = 4/5 * 60 minutes = 48 minutes